OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

Directrix (directrix):

@shannon342 Select one of these, and I'll show you how to get started. Then, you can choose a second one and factor it. Okay?

OpenStudy (anonymous):

ok i choose 8x^2-11x+3 and 5x^2 -2x-7

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

I am choosing this one --> 8x^2-11x+3

Directrix (directrix):

8x²-11x+3 is a trinomial with general expression a*x² + b*x + c. In 8x²-11x+3, a = 8, b = -11, and c = 3. To factor 8x²-11x+3 by grouping, two numbers that multiply to a*c and add to b are required. So, the first task is to find two numbers that multiply to 8*3 and add to -11. Two such numbers are -8 and -3. Next, ....

Directrix (directrix):

... break apart the -11*x into -8*x and -3*x. That gives 8x²-11x+3 = 8x² -8x -3x+3.

Directrix (directrix):

8x² -8x -3x+3 = 8x*(x - 1) - 3*(x - 1) --> This is the start of factoring by grouping.

Directrix (directrix):

Notice that (x - 1) is a factor of both 8x*(x - 1) and - 3*(x - 1). (x - 1)* [ 8x - 3 ] is now the factored form of 8x²-11x+3. @shannon342

OpenStudy (anonymous):

whoa thats alot yeah i see that

Directrix (directrix):

Well, the instructions said to write an explanation in sentences. You are up for the factorization of the second polynomial! @shannon342

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

was that the answr

OpenStudy (anonymous):

or do u want me to finish it?

Directrix (directrix):

8x²-11x+3 = (x - 1)*( 8x - 3 )

Directrix (directrix):

This is the one for you to factor and tell how to do it in complete sentences. So, the problem is at this moment only 50% done. Yours --> 5x^2 -2x-7 >>>Choose two trinomials from the list below to factor - You chose two. I did one, and we agreed that you would do the second.

OpenStudy (anonymous):

can we work the second one together? :)

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

This is the deal we made. >> Select one of these, and I'll show you how to get started. Then, you can choose a second one and factor it. Okay? >>>ok i choose 8x^2-11x+3 and 5x^2 -2x-7 I did mine. Now, it's time for you to start on yours. You can use mine as a pattern for yours.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

Directrix (directrix):

If you get started, then I can check what you write.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

5x^2-2x-7 a=5 b=-2 c= -7

OpenStudy (anonymous):

= to (5x-7)(x+1)?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

The factorization is correct but I don't know about the complete sentence explanation stuff.

OpenStudy (anonymous):

its fine thanks for ur help

Directrix (directrix):

Alrighty, then. You're welcome.